OpenStudy (mony01):

if d/dx sec(x) = sec(x)tan(x). What is d/dx (sec^-1(x))?

hartnn (hartnn):

you want to find d/dx (sec^-1 x) using that formula only?

hartnn (hartnn):

let y = sec^-1 x x= sec y differentiate both sides w.r.t y, what u get ?

OpenStudy (mony01):

how do i differentiate both sides?

hartnn (hartnn):

"d/dx sec(x) = sec(x)tan(x)." that came by differentiation sec x so, d/dy of sec y = ... ?

OpenStudy (mony01):

is it sec(y)tan(y)?

hartnn (hartnn):

yes!

hartnn (hartnn):

so, dx/dy = sec y tan y now use the fact that dy/dx = 1/ (dx/dy) so dy/dx = ...?

OpenStudy (mony01):

is it 1/sec y tan y

hartnn (hartnn):

yes now we just need to find sec y and tany in terms of x sec y is already = x can you try to find tan y in terms of x ?

hartnn (hartnn):

hint : sec^2 y + tan^2 y =1

OpenStudy (mony01):

is it sec^2(x)-1

hartnn (hartnn):

that would be tan^2x we need tan y tan y = \(\sqrt{\sec^2y-1}\) with sec y =x what will be tan y ?

OpenStudy (mony01):

\[\sqrt{x ^{2}-1}\]

hartnn (hartnn):

yes, so d/dx (sec^-1 x) = 1/ [x (\(\sqrt {x^2-1}\))] thats it!

OpenStudy (mony01):

yay thank you for helping me with this problem!!!

hartnn (hartnn):

welcome ^_^